# Align Your Intents: Offline Imitation Learning via Optimal Transport

### v1

@@ 4 Method
We extend the approach by Luo et al. Luo et al. (2023) to compute intrinsic rewards through an optimal transport. For two trajectories $\{s^{a}_{i}\}_{i=1}^{T_{a}}$ and $\{s^{e}_{j}\}_{j=1}^{T_{e}}$, the optimal transition matrix is given by solving the entropy-regularized OT problem:
@@
where $C_{ij}=c(s^{a}_{i},s^{a}_{i+k},s^{e}_{j},s^{e}_{j+k})$ is some cost function with parameter $k>0$ and $P$ has the following marginal distributions $\forall i,j$: $\sum_{j=1}^{T_{e}}P_{ij}=1/T_{a},\quad\sum_{i=1}^{T_{a}}P_{ij}=1/T_{e}$. By finding the optimal matrix $P^{*}$, we can determine the reward for each state of the trajectory $\{s^{a}_{i}\}_{i=1}^{T_{a}}$ as
$r(s_{i}^{a})=A\exp\left(-\tau T_{a}\sum_{j=j_{1}}^{T_{e}}P^{*}_{ij}C_{ij}\right),$ (6)
@@
the exponent function with the hyperparameters $A$ and $\tau$ serves as an additional scaling factor to diminish the impact of the states with a large total cost. The negative sign ensures that, in the process of maximizing the sum of rewards, we are effectively minimizing the optimal transport (OT) distance.
@@
Parameters: $A$, $\tau$ – scaling, $k$ – shift in intents pair, $\varepsilon$ – OT entropy coefficient
1:  Let rewards $=$ List().
2:  Train $\psi$ intents mapping by ICVF procedure.
@@
6:        Compute cost matrix $C$ by expression (8);
7:        Let $j_{1}={\arg\min}_{j}C_{1j}$;
@@
8:        Solve OT($\left(\{s^{a}_{i}\}_{i=1}^{T_{a}},\,\{s^{e}_{j}\}_{j=j_{1}}^{T_{e}}\right)$ and obtain $P^{*}$;
9:        Compute rewards $\{r(s^{a}_{i})\}_{i=1}^{T_{a}}$ by expression (6);
10:        Append $\{r(s^{a}_{i})\}_{i=1}^{T_{a}}$ to $R$.
11:     end for
12:     For each $i$ append $r_{i}=\min_{t}R_{ti}$ to rewards.
13:  end for
14:  return rewards
@@
We proceed to estimate the transition matrix $P^{*}$ between each trajectory of the agent and the expert’s demonstrations (which may consist of one or more trajectories). For enhanced alignment, we selectively consider the tail of the expert’s trajectory $j_{1},\ldots,T_{e}$, focusing on the nearest first states to the agent’s starting position according to the cost matrix $C$. The maximum reward is selected across different trajectories when the expert provides multiple trajectories. The aggregated rewards for each trajectory of the agent are then incorporated into the offline dataset for subsequent RL training, with the policy trained by the IQL offline RL algorithm Kostrikov et al. (2021). A comprehensive recipe for the rewards computation procedure is shown in Algorithm 1.
@@
In contrast to the OTR method by Luo et al. Luo et al. (2023) that computes distances directly between states, our approach measures differences between intents. Intents are such state representations that account for the temporal path length between the states in the environment. Specifically, we use the following cost matrix:
$C_{ij}=\|\psi(s^{a}_{i})-\psi(s^{e}_{j})\|^{2}+\|\psi(s^{a}_{i+k})-\psi(s^{e}_{j+k})\|^{2}.$ (8)
The inclusion of the second term in the cost is necessary for an ordered comparison of the trajectories. Hence, during training, we enforce the distribution of the agent’s intent pairs to converge to the empirical measure of the intent pairs of the expert.
Dataset
IQ-Learn SQIL ORIL SMODICE AILOT
@@
Dataset
IQL OTR CLUE AILOT
@@
Table 1: Normalized scores (mean $\pm$ standard deviation) of AILOT on MuJoco locomotion tasks, compared to baselines. For OTR, CLUE, and AILOT, the results are given for $K=1$ number of expert trajectories. The highest scores are highlighted.
@@
5 Experiments
In the current section, we demonstrate performance of AILOT on several benchmarking tasks, make ablation study on varying number of provided expert demos and provide implementation details.
First, we empirically show the ability of proposed method to efficiently utilize expert demonstrations on sparse-rewards tasks (such as Antmaze and Adroit environments) and improve learning ability of offline RL algorithms by providing geometrically aware dense reward signal to agent’s transitions. Second, we evaluate how well AILOT performs in offline Imitation Learning (IL) setting, showing improved performance upon state-of-the art offline IL algorithms on MuJoco locomotion tasks. Moreover, for imitation learning, we include additional experiments with custom locomotion behaviors, e.g expert hopper performing backflip, when agent dataset consists of only completely random behaviors.
5.1 Implementation
Because our method provides intrinsic dense reward relabelling, any offline RL algorithm can be used afterwards to learn decision making policy, closely resembling expert behaviors. In all our experitiments we endow AILOT with Implicit Q-Learning (IQL), which is simple and robust offline RL algorithm.
We implement AILOT in JAX Bradbury et al. (2018) and use official implementations of both ICVF  [^1] https://github.com/dibyaghosh/icvf_release and IQL  [^2] https://github.com/ikostrikov/implicit_q_learning. Our code is written using Equinox library Kidger and Garcia (2021). In order to compute optimal couplings, we utilize efficient implementation of Sinkhorn algorithm Cuturi (2013) from OTT-JAX library Cuturi et al. (2022). We found that $200$ Sinkhorn solver iterations are enough to find optimal mapping and set $\epsilon=0.001$ as entropy regularization parameter. All hyperparameters for IQL on D4RL benchmarks are set to those recommended in original paper to ensure reproducibility. We use default latent dimension of $256$ for ICVF embeddings pretraining across all benchmarks and set other hyperparameters as in original paper. In order to maintain rewards in a reasonable range, they are scaled by exponential function as written in Equation 6 with $A=5,\tau=0.5$ for MuJoco and AntMaze tasks and $A=5,\tau=10$ for Adroit. We found that lookahead parameter in Equation 8 works best for $k=2$. Results are evaluated across 10 random seeds and 10 evaluation episodes for each seed in order to be consistent with previous works.
Figure 2: Expert demonstrator Hopper performs a backflip.
Figure 3: Hopper agent successfully performs imitation of backlip from the observations via AILOT.
Figure 4: Expert demonstrator HalfCheetah performs an upward standing.
Figure 5: HalfCheetah agent successfully performs imitation of standing upwards from the observations via AILOT.
@@ 5.1 Implementation
Figure 6: The squared norm of states and intents differences, depending on the total steps count between the states in the same trajectory. The average squared norm of intents differences has a near-linear dependence on the steps count. The squared norm in the state space is not a monotone function, which is less efficient for training an imitating agent.
5.2 Baselines

### v2

@@ 4 Method
Opposite to Luo et al. (Luo et al., 2023), we change perspective of aligning occupancy measures from the state space into the alignment of intents in the metric-aware latent space by performing isometric transformation by terms of temporally preserving function (Section 3.3). First, we briefly outline how the optimal transportation is computed in the state space.
For two trajectories $\{s^{a}_{i}\}_{i=1}^{T_{a}}$ and $\{s^{e}_{j}\}_{j=1}^{T_{e}}$, the optimal transition matrix is obtained by solving the entropy-regularized (Sinkhorn) (Cuturi, 2013) OT problem:
@@
where $C_{ij}=c(s^{a}_{i},s^{a}_{\min(i+k,T_{a})},s^{e}_{j},s^{e}_{\min(j+k,T_{e})})$ is some cost function with the parameter $k>0$ and $P$ has the following marginal distributions $\forall i,j$: $\sum_{j=1}^{T_{e}}P_{ij}=1/T_{a},\quad\sum_{i=1}^{T_{a}}P_{ij}=1/T_{e}$, which we take to be uniform across both trajectories. By finding the optimal matrix $P^{*}$, we can determine the reward for each state of the trajectory $\{s^{a}_{i}\}_{i=1}^{T_{a}}$ as
$r(s_{i}^{a})=\alpha\exp\left(-\tau T_{a}\sum_{j=j_{1}}^{T_{e}}P^{*}_{ij}C_{ij}\right),$ (8)
@@
and the exponent function with a scalar hyperparameters $\alpha$ and $\tau$ serves as an additional scaling factor to diminish the impact of the states with a large total cost. The negative sign ensures that, in the process of maximizing the sum of rewards, we are effectively minimizing the optimal transport (OT) distance.
@@
Parameters: $\alpha$, $\tau$ – scaling coefficients, $k$ – intents pair shift, $\varepsilon$ – OT entropy coefficient
1:  Train $\psi$ intents mapping by ICVF procedure (7).
2:  Let rewards $=$ List();
@@
6:        for  $1\leq i\leq T_{a}$ and $1\leq j\leq T_{e}$ do
7:           Set $i\_next=\min(i+k,T_{a})$ and $j\_next=\min(j+k,T_{e})$;
8:           Compute cost $C_{ij}=\|\psi(s^{a}_{i})-\psi(s^{e}_{j})\|^{2}+\|\psi(s^{a}_{i\_next})-\psi(s^{e}_{j\_next})\|^{2}$;
9:        end for
10:        Let $j_{1}={\arg\min}_{j}C_{1j}$;
@@
11:        Solve $\text{OT}_{\varepsilon}\left(\{s^{a}_{i}\}_{i=1}^{T_{a}},\,\{s^{e}_{j}\}_{j=j_{1}}^{T_{e}}\right)$ and obtain transport matrix $P^{*}$;
12:        For $1\leq i\leq T_{a}$ compute rewards $r(s_{i}^{a})=\alpha\exp\left(-\tau T_{a}\sum_{j=j_{1}}^{T_{e}}P^{*}_{ij}C_{ij}\right)$;
13:        Append $\{r(s^{a}_{i})\}_{i=1}^{T_{a}}$ to R.
14:     end for
15:     For each $i$ append $r_{i}=\min_{t}\texttt{R}_{ti}$ to rewards (the min by expert trajectories).
16:  end for
17:  return rewards
@@
We proceed to estimate the transition matrix $P^{*}$ between each trajectory of the agent and the expert’s demonstrations (which may consist of one or more trajectories) as described in Section 3.2. For enhanced alignment, we selectively consider the tail of the expert’s trajectory $j_{1},\ldots,T_{e}$, focusing on the nearest first states to the agent’s starting position according to the cost matrix $C$. The maximum reward is selected across different trajectories when the expert provides multiple trajectories. The aggregated rewards for each trajectory of the agent are then incorporated into the offline dataset for subsequent RL training, with the policy trained by the IQL offline RL algorithm (Kostrikov et al., 2021). We use ICVF representations only in rewards definition and employ IQL without any modifications. A comprehensive recipe for the rewards computation is shown in Algorithm 1.
@@
Unlike the OTR method (Luo et al., 2023) which computes distances directly between the states, our approach measures differences between intentions $\psi(s)$. Specifically, we propose to incorporate the following cost matrix:
$C_{ij}=\|\psi(s^{a}_{i})-\psi(s^{e}_{j})\|^{2}+\|\psi(s^{a}_{\min(i+k,T_{a})})-\psi(s^{e}_{\min(j+k,T_{e})})\|^{2}.$ (10)
The second term in the cost function is necessary for an ordered comparison of the trajectories. In imitation learning, we always compare distributions of pairs of the expert and the agent states. This is so because we want to be not only in the same state, but we also want to act in a manner similar to the expert. Hence, during training, we enforce the distribution of the agent’s intent pairs to converge to the empirical measure of the intent pairs of the expert. The motivation behind the extraction of the behaviors comes from the observation that the time-step differences in the state space provide erroneous values, not capturing temporal dependencies between the dataset states, which we show in Figure 3, and thus, the geometry of the environment is not represented properly, making it hard for the optimal transport to scale in high-dimensional problems.
5 Experiments
In this Section, we demonstrate the performance of AILOT on several benchmark tasks, report an ablation study on varying the number of provided expert demonstrations, and provide the implementation details. First, we empirically show the ability of proposed method to efficiently utilize expert demonstrations in sparse-reward tasks (such as Antmaze and Adroit environments) and improve learning ability of Offline RL algorithms by providing geometrically aware dense intrinsic reward signal to agent’s transitions. Second, we evaluate AILOT in the offline Imitation Learning (IL) setting, outperforming the state-of-the art offline IL algorithms on MuJoco locomotion tasks. Also, we include additional experiments with custom behaviors, e.g., the expert hopper performing a backflip, when the agent dataset consists of completely random behaviors.
5.1 Implementation Details
Dense reward relabelling by AILOT is completely decoupled from the offline policy training. In all our experiments, we endow AILOT with Implicit Q-Learning (IQL) (Kostrikov et al., 2021), which is a simple and a robust offline RL algorithm. In our additional experiments, we also test AILOT+IQL against Diffusion-QL (Wang et al., 2022), which is a recent state-of-the-art approach for offline RL.
We implement AILOT in JAX (Bradbury et al., 2018) and use the official IQL implementation [^1] https://github.com/ikostrikov/implicit_q_learning. Our code is written using Equinox library (Kidger & Garcia, 2021). To compute optimal transport, we use implementation of Sinkhorn algorithm (Cuturi, 2013) from OTT-JAX library (Cuturi et al., 2022). Additional details on chosen hyperparameters and settings are provided in the Appendix A.2.
Figure 2: Top: Sample trajectory from the agent dataset, showcasing backflip task. Bottom: Hopper agent successfully performs imitation of backflip from the observations via AILOT. Refer to the Supplementary material for the animations.
5.2 Runtime
All of the experiments were conducted on a single RTX 3090. The runtime overhead of the application of optimal transport is not greater than extra $\sim 10$ minutes (compared to the offline RL algorithm training time). Thus, the total runtime of AILOT+IQL equals $\sim$ 25 minutes. Finding optimal matrix $P^{*}$ and correspondent internal rewards takes several hundred iterations of Sinkhorn solver, which is relatively fast. Moreover, because the optimal transport is computed in a fixed latent space of intents, the overhead stays constant across broad range of tasks regardless of the state dimensionality. For high-dimensional tasks (e.g., learning from pixels), AILOT is more suitable than OTR due to performing OT in the latent space, accelerating the overall relabelling.
5.3 Baselines
AILOT is compared against the following state-of-the-art algorithms: IQL (Kostrikov et al., 2021), OTR (Luo et al., 2023), CLUE (Liu et al., 2023), IQ-Learn (Garg et al., 2021), Diffusion-QL (Wang et al., 2022), SQIL (Reddy et al., 2019), ORIL (Zolna et al., 2020), and SMODICE (Ma et al., 2022a). The details are given in Appendix B.
Dataset IQ-Learn SQIL ORIL SMODICE AILOT
@@ 5.3 Baselines
Dataset IQL OTR CLUE AILOT
@@
Table 1: Normalized scores (mean $\pm$ standard deviation) of AILOT on MuJoco locomotion tasks, compared to baselines. The upper sub-table includes methods for imitation learning without any kind of optimal transport. The lower sub-table shows results for the conceptually close approaches – oracle IQL, OTR, and CLUE. For these methods the results are given for $K=1$ number of expert trajectories. The highest scores are highlighted in green.
@@
Table 2: Normalized scores (mean $\pm$ standard deviation) of AILOT on sparse-reward environments. We compare different dense relabelling methods (OTR, CLUE) and show that AILOT outperforms those approaches and accelerates learning of offline IQL method.
@@
5.4 Results & Experiments
We present the results for the following offline RL settings: (1) the imitation setting, where the goal is to mimic the behavior as closely as possible given several expert demonstrations in the form of trajectories; (2) the sparsified reward tasks, where the reward of one is given only when the agent reaches the goal state. We show that the rewards obtained through AILOT relabelling in both settings are descriptive enough to recover the demonstrated policy. We evaluate proposed method on common D4RL (Fu et al., 2020) benchmark and consider following domains: Gym, Adroid and Antmaze. Additional details on the environments are included in the Appendix F.
Offline Imitation Learning.
We compare AILOT performance on IL tasks in the offline setting. D4RL MuJoco locomotion datasets are used as the main source of offline data with the original reward signal and the action labels discarded, and the expert trajectories chosen for each task as the best episodes achieving maximal return. Results for D4RL locomotion tasks are presented in Table 1. AILOT achieves the best performance in $7$ out of $9$ benchmarks, when compared to OTR and CLUE. Note that, unlike the CLUE method, our algorithm does not rely on the action labels. Empirically, we confirm that the optimal geometrically aware map between the expert and the agent trajectories in the intents latent space gives reasonable guidance for the agent to learn properly. Since representation $\psi(s)$ was pretrained by general objective Eq.5 for all possible $z$, we can extract it to obtain semantic behavior of policy, which produced high return expert trajectories. We chose Euclidean distance as a measure of similarity in the intent latent space as in Eq.10 and perform exponential scaling of rewards according to Eq.8 in order to maintain them in the appropriate range.
Also, we include additional imitation results when a custom demonstrations are provided, as in Figure 2 and Figure 5, where the Hopper expert makes a backward flip and the HalfCheetah stands upwards respectively. For those tasks, agent’s dataset initially consists of only the random behaviors and agent learns to recover the desired behavior through the AILOT dense relabelling.
Sparse-Reward Offline RL Tasks.
Next, we evaluate AILOT on several sparse D4RL benchmarks (AntMaze-v2 and Adroit-v2). To obtain expert trajectories, we consider only those episodes that accomplish the goal task, dismissing all the others.
Table 3 compares the performance of AILOT+IQL to OTR+IQL,  CLUE+IQL, when only a single demonstration trajectory is available, and to the original IQL. We employed IQL, which acts as a baseline, without any modifications. Given that OTR + IQL always performs better than IQL, we decided to compare directly with OTR, which showcases the importance of performing optimal transport alignment under distance preserving mapping. We observe that AILOT outperforms the current state-of-the art results, with the hardest antmaze-large-diverse task showing the most remarkable margin. This proves the ability of AILOT to employ the expressive representations from general pretrained value function (5) through the optimal transport for functional learning in the sparse tasks as well. In door-cloned-v0 and hammer-cloned-v0 tasks IQL performs best with sparse rewards. The inferior results in those two tasks stem from the fact that the cloned version was obtained from collecting the data from trained imitation learning policy on a mix of human and the expert data. Human data contains many ambiguous trajectories, from which it is hard to extract valuable intents (this is the case for both the door and the hammer-cloned tasks).
5.5 Ablation Study
Varying the number of expert trajectories.
We investigate whether performance of learned behavior tends to improve with increased number of provided expert trajectories. Table 7 in the Appendix shows overall performance for varying number of expert trajectories for $K=1$ to $K=5$ across OTR and AILOT. However, we observe that performance across both algorithms improves slightly. We make comparison with OTR here because it’s the most similar to ours (it also uses OT for reward labelling and IQL for RL problem solution). Still, AILOT achieves better normalized scores than OTR, thus proving that alignment of intents in geometry-aware space improves intrinsic rewards labelling in comparison with similar rewards but with pairwise distances between original states as done in OTR.
@@ 5.5 Ablation Study
Figure 3: The squared norm of states and intents differences, depending on the total steps count between the states in the same trajectory. The intents differences $||\psi(s_{t+k})-\psi(s_{t})||^{2}$ has a near-linear dependence on the steps count. The squared norm in the state space is not a monotone function, which is less efficient for training an imitating agent, since it completely ignores global geometric dependencies between states in the dataset.
Intents distance dependence on the steps count.
By learning temporally preserving value function (i.e., similar states get mapped to temporally closed points in the intents space), the underlying space becomes more structured. Because the optimal transport takes the geometry of space into account, the task of finding the alignment becomes simpler. In the Method section, we have already mentioned that $\mathbb{E}_{t}\|\psi(s_{t+k})-\psi(s_{t})\|^{2}$ is near-linearly dependent on the steps count between the states ($k$). This important property plays a constructive role in the cost matrix in the OT problem (10), and ultimately, gives a good estimate of the distance between the trajectories. Empirical evidence of the near-linear dependence is presented in Figure 3. On this figures we show that the pairwise distances between the original states $\mathbb{E}_{t}\|s_{t+k}-s_{t}\|^{2}$, which is what used in previous OT methods, have no such feature and their direct use as a cost function is less preferable, since they discarding global temporal geometry. Additionally, we compare AILOT to OTR with the same $\alpha$ and $\beta$ hyperparameters in the Appendix, showcasing that performance gains come from performing OT in temporally grounded latent space.
6 Discussion
In our work, we introduced AILOT – a new non state-occupancy estimation method for extracting expert’s behavior in terms of its intentions and guiding the learning agent to them through an intrinsic reward. We empirically show that it surpasses the state-of-the art results both in the sparse-reward RL tasks and in the offline imitation learning setting. AILOT can mimic the expert behaviour without knowing its action labels, and without the ground truth rewards. Moreover, we show that the intrinsic rewards, distilled by AILOT, could be used to efficiently boost the performance of other offline RL algorithms, thanks to the proper alignment to the expert intentions via the optimal transport.
Limitations
of AILOT include assumption on sufficient access to a large number of unlabeled trajectories of some acceptable quality. In our work, we have focused on the expert behaviours, typically considered in the offline RL publications: popular synthetic environments with some comprehensible expert movements and, consequently, some sufficiently ‘intuitive’ intentions, which can be easily extracted from the provided datasets.
Another limitation follows from the multi-modality of intents, because the expert can have several goals or performs a vague action. The imitation efficiency could, of course, drop, because the expert intents may no longer be transparent to the agent. While such a trait would be on par with the way humans learn a certain skill by observing an adept, weighing the hierarchy of multiple possible intents in AILOT could prove useful to further regularize the learning dynamics in such uncertain scenarios and is an interesting direction for further research. Other direction of future work is to venture into the cross-domain imitation. Based on the results observed here, it should be possible to generalize AILOT to handle the transition shift between the expert and the agent in the presence of larger mismatches pertinent to the different domains.
In conclusion, the development of AILOT sets a robust benchmark for future generalizations, enhancing ongoing research in crafting generalist knowledge distillation agents from the offline data.
7 Reproducibility Statement
We included source code in the supplementary material. All of the hyperparameters used, along with additional dataset preprocessing step, are discussed in Appendix A.2. In our experiments, we used open-source D4RL datasets (Fu et al., 2020). In order to ensure reproducibility and account for randomness, we repeated our experiments over 10 random seeds.
Appendix A Appendix
Figure 4: Principal diagram of AILOT approach for imitation learning.
A.1 Value function convergence
Proposition 1
Consider all possible pairs of current and goal states ($s,s_{+}$) with $z=\psi(s_{+})$. Assume that functions $\phi(s)$, $\psi(s_{+})$, $T(\psi(s_{+}))$ (decomposition of $V(s,s_{+},z)$ from eq. 5) and expression
$\left\|\phi^{T}(s)\frac{\partial T(\psi_{0})}{\partial\psi}\psi(s_{+})\right\|$ (11)
are upper-bounded by some constant value. Here $\psi_{0}$ is some intermediate value between $\psi(s)$ and $\psi(s_{+})$. Then from convergence of intent embedding $\psi(s)\to\psi(s_{+})$ (w.r.t Euclidean norm) it follows
$V(s,s_{+},z)\to 0.$
Proof: Define $V(s,s_{+})=V(s,s_{+},\psi(s_{+}))$ and $\delta=\|\psi(s)-\psi(s_{+})\|.$ The goal is to find an upper bound $V(s,s_{+})$. Since $V(s,s)=0$, $-V(s,s_{+})$ can be rewritten as
$-V(s,s_{+})=-V(s,s_{+})+V(s,s).$ (12)
Take by definition $V(s,s_{+})=\phi^{T}(s)T(\psi(s_{+}))\psi(s_{+})$ and obtain
$-V(s,s_{+})$ $=\phi^{T}(s)T(\psi(s))\psi(s)-\phi^{T}(s)T(\psi(s_{+}))\psi(s_{+})$ (13)
$=\phi^{T}(s)T(\psi(s))[\psi(s)-\psi(s_{+})]+\phi^{T}(s)[T(\psi(s))-T(\psi(s_{+}))]\psi(s_{+})$ (14)
$\leq\|\phi^{T}(s)T(\psi(s))\|\delta+\left\|\phi^{T}(s)\frac{\partial T(\psi_{0})}{\partial\psi}\psi(s_{+})\right\|\delta.$ (15)
Since the norms in the expression above are upper-bounded then from $\delta\to 0$ follows $V(s,s_{+})\to 0$.
A.2 Hyperparameters
We found that $n=200$ Sinkhorn solver iterations are enough to find optimal coupling and set $\epsilon=0.001$ as entropy regularization parameter. All hyperparameters for IQL on D4RL benchmarks are set to those recommended in original paper to ensure reproducibility. We use default latent dimension of $d=256$ for general value function from Eq. 5 embeddings pretraining across all benchmarks and set other hyperparameters as in original paper. In order to maintain rewards in a reasonable range, they are scaled by exponential function as written in Equation 8 with $\alpha=5,\tau=0.5$ for MuJoco and AntMaze tasks and $\alpha=5,\tau=10$ for Adroit. We found that lookahead parameter in Eq. 10 works best for $k=2$. Results are evaluated across 10 random seeds and 10 evaluation episodes for each seed in order to be consistent with previous works.
It should be noted that IQL incorporates its own rewards rescaling function within the dataset. We apply similar technique, using the reward scaling factor of $\frac{1000}{\mbox{max\_return}-\mbox{min\_return}}.$
For AntMaze-v2 tasks, we subtract 1 from the rewards outputted by AILOT. All other parameters for IQL for other tasks are kept intact. Full list of crucial parameters are presented in Tables 4 and 5. For AntMaze tasks, the parameters are the same as for Mujoco, except for the expectile in IQL, which we set to $0.9$. Pretraining of general value function is executed for $250$k steps, which we found to be enough to provide reasonable distance estimation, which coincides with original ICVF implementation.
Algorithm Hyperparameter Value
IQL Temperature 6
Expectile 0.7
Hidden layers (256, 256)
Optimizer Adam
Critic learning rate $3e^{-4}$
Value learning rate $3e^{-4}$
Policy learning rate $3e^{-4}$
AILOT $\tau$ 0.5
$\alpha$ 5
Sinkhorn $\epsilon$ 0.001
k 2
Table 4: Hyperparameters for MuJoco tasks.
Algorithm Hyperparameter Value
IQL Temperature 5
Expectile 0.7
AILOT $\tau$ 10
$\alpha$ 5
Sinkhorn $\epsilon$ 0.001
k 2
Table 5: Hyperparameters for Adroit tasks. IQL parameters are the same as for the locomotion tasks.
Appendix B Details about the Baseline Models